OpenStudy (zxcvbnm1216):

How is the electrical force related to the magnitude of the charge on the one of the objects?

ganeshie8 (ganeshie8):

Hint : Coulombs law

OpenStudy (zxcvbnm1216):

@ganeshie8 what do you mean? I am supposed to know an equation or something?

ganeshie8 (ganeshie8):

Yes. Lookup coulombs law

OpenStudy (aakashtomar):

(Kq1q2)/r^2 = F where K is reciprocal of permitivity ( means it is constant ) and q1; q2 are charges. r is the separation between them.

OpenStudy (agent0smith):

If you've been given this question, yes you're probably expected to be aware of coulombs law... look at the numerator of the formula.

OpenStudy (nincompoop):

Coulumb's Law: force between two point charges: \(F = \dfrac{1}{4\pi \epsilon_0} \dfrac{\lvert q_1q_2 \lvert}{r^2} \) where: \(\epsilon_0 = 8.854 \times 10^{-12} C/N \cdot m^2 \rightarrow \dfrac{1}{4\pi \epsilon_0} = k = 8.988 \times 10^{9} N \cdot m^2/C^2 \)